In McDonald’s restaurants with automated kiosks to take orders, the kiosks can take 2 orders per...

  1. In McDonald’s restaurants with automated kiosks to take orders, the kiosks can take 2 orders per minute while employees can take 3 orders per minute. The rental rate on a kiosk is $6 per hour while the average compensation costs for an employee is $18 per hour. Should McDonald’s install more kiosks?
Answer #1


The kiosks can take 2 orders per minute.

So,

MPK = 2

The rental rate of kiosks is $6 per hour.

PK = $6

An employee can take 3 orders per minute.

So,

MPL = 3

The wage of employee is $18 per hour.

PL = $18

McDonald will minimize cost when the following condition is fulfilled -

MPK/PK = MPL/PL

MPK/PK = 2/6 = 0.33

MPL/PL = 3/18 = 0.17

So,

MPK/PK > MPL/PL

This means that McDonald should increase the quantity of kiosks and reduce the number of workers to attain the cost minimizing combination.

Thus,

McDonald should install more kiosks.